Find the slope and y-intercept for the following equation by writing it in the form y = mx + b.
2 x + 5 y = 20

Found 2 solutions by mananth, blwinbbbles:
Answer by mananth(16946)

2x+5y=20
-2x
2x-2x+5y=20-2x
5y=-2x+20
/5
y=(-2/5)x+4
slope = (-2/5)
plug x =0 for y intercept
y=4

Answer by blwinbbbles(106)

1) y = mx + b
m = slope b = the y intercept
2) 2x + 5y = 20
First rearrange equation 2 to the form of equation 1
So: subtract 2x from both sides
5y = -2x + 20
Now divide by 5 all the way through
y = -2/5(x) + 4
The slope is -2/5 and the intercept is 4

hope this helps